Calculate the wavelength.
A laser shines on a pair of vertical slits. The horizontal distance L1 between the laser and the ruler is 12.3 m. The distance L2 between the laser and the slits is 0.511 m. The distance d is 0.340 m. Intensity4 2 Id 0
Answer #1

Length, L = 12.3 - 0.511 = 11.789 m

Wavelength, lambda = dy/L

Lambda = (0.39 x 10^-3 x 0.032)/11.789 = 1.06 x 10^-6 m

Comment in case any doubt please rate my answer...